Two soccer players start from rest, 49 m apart. They run directly toward each other, both players accelerating. The first player’s acceleration has a magnitude of 0.40 m/s2. The second player’s acceleration has a magnitude of 0.50 m/s2. (a) How much time passes before the players collide? (b) At the instant they collide, how far has the first player run?


1
Expert's answer
2021-10-07T08:48:14-0400

(a) Let "L=49\ m" is the distance between two soccer players, "d" is the distance that the first player runs before the players collide. Then, we can write the displacements of each soccer player:


"d_1=\dfrac{1}{2}a_1t^2=d,""d_2=\dfrac{1}{2}a_2t^2=L-d."

Let's substitute "d" from the first equation into the second one and find the time that passes before the players collide:


"\dfrac{1}{2}a_2t^2=L-\dfrac{1}{2}a_1t^2,""\dfrac{1}{2}t^2(a_1+a_2)=L,""t=\sqrt{\dfrac{2L}{(a_1+a_2)}}=\sqrt{\dfrac{2\cdot49\ m}{(0.4\ \dfrac{m}{s^2}+0.5\ \dfrac{m}{s^2})}}=10.4\ s."

(b) We can find the distance that the first player runs before the players collide as follows:


"d=\dfrac{1}{2}a_1t^2=\dfrac{1}{2}\cdot0.4\ \dfrac{m}{s^2}\cdot(10.4\ s)^2=21.6\ m."